3. Consider a consumer who wants to consume only two commodities and has an income of $100. Assume the price of good 1 is $10 per unit and the price of good 2 is $20 per unit. Now, inflation causes the price of good 1 to increase to $20 per unit, while the price of good 2 increases to $25 per unit. On the other hand, the consumer also gets a raise of $100 (so her new income is $200). Is she better off or worse off?


1
Expert's answer
2021-08-03T12:55:48-0400

Originally, the consumer’s budget constraint is

"10\times x1+20\times x2 \le100"

The budget line has horizontal intercept

"\frac{m}{p1}=\frac{100}{10}=10"

and vertical intercept

"\frac{m}{p2}=\frac{100}{20}=5"

After the change, her budget constraint becomes

"20\times x1+25\times x2 \le200"

and so the new budget line has horizontal intercept

"\frac{m}{p1}=\frac{200}{20}=10"

 (the same as before) and vertical intercept

"\frac{m}{p2}=\frac{200}{25}=8" (higher than before)

and so the consumer’s budget set has grown: she can now afford bundles she previously could not, and she can still afford all bundles she previously could. Therefore she is better off.